OpenStudy (firejay5):

Explain why log(-100) is not defined.

OpenStudy (perl):

lets go over how we 'read' logs. What does log(100) mean?

OpenStudy (abb0t):

Have you tried looking at the graph of log(x)

OpenStudy (anonymous):

\[\log(-100)=y\iff 10^y=-100\] try solving and see what happens

OpenStudy (abb0t):

Assuming you're working with \(\sf \mathbb{R}\)

OpenStudy (firejay5):

@perl @abb0t @satellite73 the answer to my question was simply, because the inner expression of a log must be postive

OpenStudy (abb0t):

Ok.

OpenStudy (abb0t):

All you needed to do was observe the graph